# Latent Factor Analysis in Short Panels

### v2

@@ B Proofs of Propositions 1-5
We apply Lemma 3 to sequences $\tilde{c}_{l}(\lambda)$ and $\gamma_{l}$. We detail the case $df\geq 3$ (for $df=2$ the analysis is simpler). The separation condition $\frac{1}{2}\sum_{j=1}^{df}\rho_{j}^{i}\left[\rho_{j}+\frac{i+1}{\nu_{df}}(1-\nu_{j})\lambda_{j}^{2}\right]\geq\frac{df-1}{2}$, for $i=0$, yields $\lambda_{1}^{2}+\sum_{j=2}^{df}(1-\nu_{j})\lambda_{j}^{2}\geq\nu_{df}\left(df-2-\sum_{j=2}^{df-1}\rho_{j}\right)$, and, for $i\geq 1$, it yields $\sum_{j=2}^{df-1}\rho_{j}^{i}(1-\nu_{j})\lambda_{j}^{2}+(1-\nu_{df})\lambda_{df}^{2}\geq\frac{\nu_{df}}{i+1}\left(df-2-\sum_{j=2}^{df-1}\rho_{j}^{i+1}\right)$. Inequalities (8) follow.
C Monte Carlo experiments
This appendix gives a Monte Carlo assessment of size and power and selection procedure for the number of factors for the LR test under non-gausian errors. Let us start with a description of the DGP we use in our simulations. In the DGP, the betas are $\beta_{i}~{}\overset{i.i.d.}{\sim}~{}N(0,I_{k})$, with $k=3$, and the matrix of factor values is $F=V_{\varepsilon}^{1/2}U\Gamma^{1/2}$, where $U=\tilde{F}(\tilde{F}^{\prime}\tilde{F})^{-1/2}$ and $vec(\tilde{F})\sim N(0,I_{Tk})$. We generate the diagonal elements of $V_{\varepsilon}=diag(h_{1},...,h_{T})$ through a common time-varying component in idiosyncratic volatilities (Renault, Van Der Heijden and Werker (2023)) via the ARCH $h_{t}=0.6+0.5h_{t-1}z_{t-1}^{2}$, with $z_{t}\sim IIN(0,1)$. This common component induces a deviation from spherical errors. The diagonal matrix $\Gamma=Tdiag(3,2,n^{-\bar{\kappa}})$ yields $\frac{1}{T}F^{\prime}V_{\varepsilon}^{-1}F=diag(3,2,n^{-\bar{\kappa}})$, i.e., the "signal-to-noise" ratios equal $3$, $2$ and $n^{-\bar{\kappa}}$ for the three factors. We take $\bar{\kappa}=\infty$ to study the size of $LR(2)$. To study the power of $LR(2)$, we take $\bar{\kappa}=0$ to get a global alternative and $\bar{\kappa}=1/2$ to get a local alternative (weak factor). We generate the idiosyncratic errors by $\varepsilon_{i,t}=h_{t}^{1/2}h_{i,t}^{1/2}z_{i,t}$, where $h_{i,t}=c_{i}+\alpha_{i}h_{i,t-1}z_{i,t-1}^{2}$, with $z_{i,t}\sim IIN(0,1)$ mutually independent of $z_{t}$. We use the constraint $c_{i}=\sigma_{ii}(1-\alpha_{i})$ with uniform draws for the idiosyncratic variances $V[\varepsilon_{i,t}]=\sigma_{ii}~{}\overset{i.i.d.}{\sim}~{}U[1,4]$, so that $V[\varepsilon_{i,t}/h_{t}^{1/2}]=\frac{c_{i}}{1-\alpha_{i}}=\sigma_{ii}$. Such a setting allows for cross-sectional heterogeneity in the variances of the scaled $\varepsilon_{i,t}/h_{t}^{1/2}$. The ARCH parameters are uniform draws $\alpha_{i}~{}\overset{i.i.d.}{\sim}~{}U[0.2,0.5]$ with an upper boundary of the interval ensuring existence of fourth-order moments. We generate $5,000$ panels of returns of size $n\times T$ for each of the $100$ draws of the $T\times k$ factor matrix $F$ and common ARCH process $h_{t}$, $t=1,...,T$, in order to keep the factor values constant within repetitions, but also to study the potential heterogeneity of size and power results across different factor paths. The factor betas $\beta_{i}$, idiosyncratic variances $\sigma_{ii}$, and individual ARCH parameters $\alpha_{i}$ are the same across all repetitions in all designs of the section. We use three different cross-sectional sizes $n=500,1000,5000$, and three values of time-series dimension $T=6,12,24$. The variance matrix $\hat{\Omega}_{\bar{Z}^{*}}$ is computed using the parametric structure of Lemma 8. We get the $T-1$ estimated parameters by least squares, as detailed in OA Section E.5.3 i). The $p$-values are computed over $5,000$ draws.
We provide the size and power results in % in Table 1. Size of $LR(2)$ is close to its nominal level $5\%$, with size distortions smaller than $1\%$, except for the case $T=24$ and $n=500$. The impact of the factor values on size is small for $T$ above $6$. The labels global power and local power refer to $\bar{\kappa}=0$ and $\bar{\kappa}=1/2$, and power computation is not size adjusted. The global power is equal to 100%, while the local power ranges from 80% to 85% for $T=6$, and is equal to 100% for $T=12$ and $T=24$. The approximate constancy of local power w.r.t. $n$, for large $n$, is coherent with theory implying convergence to asymptotic local power. In the last panel of Table 1, we provide the average of the estimated number $\hat{k}_{LR}$ of factors, obtained by sequential testing with $LR(k)$, for $k=0,\dots,k_{max}$, with $k_{max}=2,7,17$ for $T=6,12,24$ (see Table 3 of OA). We follow the procedure described in Section 6.2, with size $\alpha_{n}=10/n$. If we reject for all $k=0,\dots,k_{max}$, then the estimated number of factors is set to $\hat{k}_{LR}=k_{max}+1$. For all sample sizes $T=6,12,24$, the average estimated number of factors is very close to the true number 2. We can conclude that our selection procedure for the number of factors works well in our simulations.
Figure 1: The upper panel displays the p-values for the statistic $LR(k)$ for the subperiods from January 1963 to December 2021, stopping at the smallest $k$ such that $H_{0}(k)$ is not rejected at level $\alpha_{n}=10/n_{max}$. If no such $k$ is found then p-values are displayed up to $k_{max}$. We use rolling windows of $T=20$ months moving forward by $12$ months each time. The first bar of p-values covers the whole 20 months. Other bars cover the last 12 months of the 20 months subperiod. We flag bear market phases with grey shaded vertical bars. The five lower panels display $\left(\overline{\hat{V}}_{y}\right)^{1/2}$ for total cross-sectional volatility, $\left(\overline{\hat{F}^{\prime}\hat{F}}\right)^{1/2}$ for systematic volatility, $\left(\overline{\hat{V}}_{\varepsilon}\right)^{1/2}$ for idiosyncratic volatility, as well as $\hat{R}^{2}$ and $\hat{R}^{2}$ under a single-factor model.
Figure 2: The upper and lower panels display the p-values for the RS and KP statistics for the subperiods from January 1963 to December 2021, for the rank test of the null hypothesis $H_{0,sp}(r)$ that $F^{O}$ has rank $r$ against the alternative hypothesis of rank larger than $r$, for any integer $r\leq k-1$. The empirical matrix $\hat{F}^{O}$ is computed with the time-varying portfolio weights of the Fama-French five-factor model plus momentum. We stop at the smallest $r$ such that $H_{0,sp}(r)$ is not rejected at level $\alpha_{n}=10/n$. If no such $r$ is found then p-values are displayed up to $k-1$. The red horizontal segments give $\hat{k}-1$, i.e., the estimated number of latent factors obtained from Figure 1 minus $1$. We flag bear market phases with grey shaded vertical bars, and use the same rolling windows as in Figure 1.
Size (%) Global Power (%) Local Power (%) $\hat{k}_{LR}$
$T$ 6 12 24 6 12 24 6 12 24 6 12 24
$n=500$ 6.0 5.2 6.7 100 100 100 80 100 100 2.0 2.0 2.1
(2.8) (0.3) (0.4) (0.1) (0.0) (0.0) (20.5) (0.0) (0.0) (0.1) (0.1) (0.2)
$n=1000$ 5.6 4.9 5.5 100 100 100 81 100 100 2.0 2.0 2.0
(2.3) (0.3) (0.3) (0.0) (0.0) (0.0) (21.1) (0.0) (0.0) (0.0) (0.0) (0.1)
$n=5000$ 5.3 5.0 4.9 100 100 100 85 100 100 2.0 2.0 2.0
(0.9) (0.3) (0.3) (0.0) (0.0) (0.0) (20.4) (0.0) (0.0) (0.0) (0.0) (0.1)
Table 1: For each sample size combination $(n,T)$, we provide the average size and power in % for the statistic $LR(2)$ (first three panels), and the average of the estimated number $\hat{k}_{LR}$ of factors obtained by sequential testing (last panel). Nominal size is $5\%$ for the first three panels, and $\alpha_{n}=10/n$ for the last panel. Global power refers to the global alternative $\bar{\kappa}=0$, and local power refers to the local alternative $\bar{\kappa}=0.5$. In parentheses, we report the standard deviations for size, power, and $\hat{k}_{LR}$ across $100$ different draws of the factor path.
ONLINE APPENDIX
Latent Factor Analysis in Short Panels
Alain-Philippe Fortin, Patrick Gagliardini, and Olivier Scaillet
We prove Lemmas 1-3 of the paper in Section D. We provide additional theory in Appendix E, namely the characterization of the pseudo likelihood and the PML estimator (E.1), the conditions for global identification and consistency (E.2), the asymptotic expansions for the FA estimators (E.3), the local analysis of the first-order conditions of FA estimators (E.4), the asymptotic normality of FA estimators (E.5), the definition of invariant tests (E.6), and proofs of additional lemmas (E.7). We give numerical checks of Inequalities (7) of Proposition 5 in Appendix F. Finally, we collect the maximum value of $k$ as a function of $T$ in Appendix G.
D Proofs of Lemmas 1-3
@@
Proof of Lemma 1: We have $\mathscr{Z}_{n}=\frac{1}{\sqrt{n}}(W\Sigma W^{\prime}-Tr(\Sigma)I_{T})$. Hence, $(\mathscr{Z}_{n})_{tt}=\frac{1}{\sqrt{n}}\sum_{i,j}(w_{i,t}w_{j,t}-1_{\{i=j\}})\sigma_{ij}=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}\zeta^{tt}_{m,n},$ with $\zeta^{tt}_{m,n}=\sum_{i\in I_{m}}[w^{2}_{i,t}-1]\sigma_{ii}+2\sum_{\underset{i<j}{i,j\in I_{m}}}w_{i,t}w_{j,t}\sigma_{ij},$ together with $(\mathscr{Z}_{n})_{ts}=\frac{1}{\sqrt{n}}\sum_{i,j}w_{i,t}w_{j,s}\sigma_{ij}=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}\zeta^{ts}_{m,n},t\neq s,$ with $\zeta^{ts}_{m,n}=\sum_{i\in I_{m}}w_{i,t}w_{i,s}\sigma_{ii}+\sum_{\underset{i<j}{i,j\in I_{m}}}w_{i,t}w_{j,s}\sigma_{ij}+\sum_{\underset{i>j}{i,j\in I_{m}}}w_{i,t}w_{j,s}\sigma_{ij},t\neq s,$ so that $vech(\mathscr{Z}_{n})=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}vech(\zeta_{m,n}),$ where $\zeta_{m,n}$ is the $T\times T$ matrix having element $\zeta^{ts}_{m,n}$ in position $(t,s)$. Hence, $vech(\mathscr{Z}_{n})$ is the row sum of a triangular array $\{vech(\zeta_{m,n})\}_{1\leq m\leq n}$ of independent centered random vectors. Let $\Omega_{m,n}:=V[vech(\zeta_{m,n})]$. Using Assumption 2 (a), we compute (i) $E[(\zeta^{tt}_{m,n})^{2}]=\sum_{i\in I_{m}}(E[w^{4}_{i,t}]-1)\sigma^{2}_{ii}+2\sum_{\underset{i\neq j}{i,j\in I_{m}}}\sigma_{ij}^{2}$; (ii) $E[(\zeta^{ts}_{m,n})^{2}]=\sum_{i\in I_{m}}E[w^{2}_{i,t}w^{2}_{i,s}]\sigma_{ii}^{2}+\sum_{\underset{i\neq j}{i,j\in I_{m}}}\sigma_{ij}^{2},t\neq s$; (iii) $E[\zeta^{tt}_{m,n}\zeta^{ss}_{m,n}]=\sum_{i\in I_{m}}E[w^{2}_{i,t}w^{2}_{i,s}-1]\sigma^{2}_{ii},t\neq s$; (iv) $E[\zeta^{tt}_{m,n}\zeta^{rp}_{m,n}]=\sum_{i\in I_{m}}E[w^{2}_{i,t}w_{i,r}w_{ip}]\sigma^{2}_{ii},r\neq p$; (v) $E[\zeta^{ts}_{m,n}\zeta^{rp}_{m,n}]$ $=\sum_{i\in I_{m}}E[w_{i,t}w_{i,s}w_{i,r}w_{i,p}]\sigma^{2}_{ii},t\neq s,r\neq p.$ It follows that $V[vech(\mathscr{Z}_{n})]=\frac{1}{n}\sum_{m=1}^{J_{n}}\Omega_{m,n}=D_{n}+\kappa_{n}I_{\frac{T(T+1)}{2}}=\Omega_{n}$. The eigenvalues of $D_{n}$ are bounded away from 00 under Assumption A.6 (b), because for any unit vector $\xi\in\mathbb{R}^{T(T+1)/2}$, we have $\xi^{\prime}D_{n}\xi\geq\frac{1}{n}\sum_{i=1}^{n}1_{i\in\bar{S}}\sigma_{ii}^{2}\xi^{\prime}V[vech(w_{i}w_{i}^{\prime})]\xi\geq\underline{c}\frac{1}{n}\sum_{i=1}^{n}1_{i\in\bar{S}}\sigma_{ii}^{2}\geq\underline{c}\left(1-\frac{1}{n}\sum_{i=1}^{n}(1-1_{i\in\bar{S}})\sigma_{ii}\right)^{2}\geq\underline{c}\left(1-\bar{C}\frac{1}{n}\sum_{i=1}^{n}(1-1_{i\in\bar{S}})\right)^{2}\geq\frac{\underline{c}}{4}$, for all $n$. We use the multivariate Lyapunov condition $\|\Omega^{-1/2}_{n}\|^{4}\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[\|vech(\zeta_{m,n})\|^{4}]\to 0$ to invoke a CLT. Since $\|A^{-1/2}\|^{4}\leq\frac{k^{2}}{\delta^{2}_{k}(A)}$ and $\|x\|^{4}\leq k\sum_{j=1}^{k}x^{4}_{j}$, for any $k\times k$ positive semi-definite matrix $A$ and $k\times 1$ vector $x$, it suffices to check that $\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[(\zeta^{ts}_{m,n})^{4}]\to 0,$ for all $t,s$. Besides, we can show that there exists a constant $M>0$, such that $E[(\zeta^{ts}_{m,n})^{4}]\leq Mb_{m,n}^{2(1+\delta)},$ for all $m,n,t,s$. We get $\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[(\zeta^{ts}_{m,n})^{4}]\leq M\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}b_{m,n}^{2(1+\delta)}=Mn^{2\delta}\sum_{m=1}^{J_{n}}B_{m,n}^{2(1+\delta)}=o(1)$, under Assumption 2 (d). Then, $\Omega_{n}^{-1/2}vech(\mathscr{Z}_{n})\Rightarrow N(0,I_{\frac{T(T+1)}{2}})$ by the multivariate Lyapunov CLT. Under Assumptions A.6 (a)-(c), $\Omega_{n}\rightarrow\Omega$ follows from the Slutsky theorem, and $\Omega$ is positive definite.
@@
$\psi^{(l)}(0)=\frac{(l-1)!}{2}\sum_{j=1}^{df}\nu_{j}^{l-1}\left[\nu_{j}+l(1-\nu_{j})\lambda_{j}^{2}\right],\quad l\geq 0.$ (D.1)
By using the Faa di Bruno formula for the derivatives of a composite function, we have $\frac{d^{l}}{du^{l}}e^{\psi(u)}=e^{\psi(u)}B_{l}(\psi^{\prime}(u),\psi^{{}^{\prime\prime}}(u),...,\psi^{(l)}(u))$, where $B_{l}$ is the $l$th complete exponential Bell’s polynomial (Bell (1934)). Hence, $\Psi^{(l)}(0)=B_{l}(\psi^{\prime}(0),\psi^{{}^{\prime\prime}}(0),...,\psi^{(l)}(0))$. The complete Bell’s polynomials satisfy the recurrence relation $B_{l+1}(x_{1},x_{2},...,x_{l+1})=\sum_{i=0}^{l}\binom{l}{i}B_{l-i}(x_{1},...,x_{l-i})x_{i+1}$. Thus, $\Psi^{(l+1)}(0)=\sum_{i=0}^{l}\binom{l}{i}\Psi^{(l-i)}(0)\psi^{(i+1)}(0)$. After standardization with the factorial term, and using equation (D.1), the conclusion follows.
@@
E Additional theory
E.1 Pseudo likelihood and PML estimator
The FA estimator is the PML estimator based on the Gaussian likelihood function obtained from the pseudo model $y_{i}=\mu+F\beta_{i}+\varepsilon_{i}$ with $\beta_{i}\sim N(0,I_{k})$ and $\varepsilon_{i}\sim N(0,V_{\varepsilon})$ mutually independent and i.i.d. across $i=1,...,n$. Then, $y_{i}\sim N(\mu,\Sigma(\theta))$ under this pseudo model, where $\Sigma(\theta):=FF^{\prime}+V_{\varepsilon}$ and $\theta:=(vec(F)^{\prime},diag(V_{\varepsilon})^{\prime})^{\prime}\in\mathbb{R}^{r}$ with $r=(k+1)T$. It yields the pseudo log-likelihood function $\hat{L}(\theta,\mu)=-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2n}\sum_{i=1}^{n}(y_{i}-\mu)^{\prime}\Sigma(\theta)^{-1}(y_{i}-\mu)=-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2}Tr\left(\hat{V}_{y}\Sigma(\theta)^{-1}\right)-\frac{1}{2}(\bar{y}-\mu)^{\prime}\Sigma(\theta)^{-1}(\bar{y}-\mu),$ up to constants, where $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i}$ and $\hat{V}_{y}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\prime}$. We concentrate out parameter $\mu$ to get its estimator $\hat{\mu}=\bar{y}$. Then, estimator $\hat{\theta}=(vec(\hat{F})^{\prime},diag(\hat{V}_{\varepsilon})^{\prime})^{\prime}$ is defined by the maximization of
$\hat{L}(\theta):=-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2}Tr\left(\hat{V}_{y}\Sigma(\theta)^{-1}\right),$ (E.1)

### v3

@@ A Proofs
We apply Lemma 4 to sequences $\tilde{c}_{l}(\lambda)$ and $\bar{\gamma}_{l}$. We detail the case $df\geq 3$ (for $df=2$ the analysis is simpler). The separation condition $\frac{1}{2}\sum_{j=1}^{df}\rho_{j}^{i}\left[\rho_{j}+\frac{i+1}{\nu_{df}}(1-\nu_{j})\lambda_{j}^{2}\right]\geq\frac{df-1}{2}$, for $i=0$, yields $\lambda_{1}^{2}+\sum_{j=2}^{df}(1-\nu_{j})\lambda_{j}^{2}\geq\nu_{df}\left(df-2-\sum_{j=2}^{df-1}\rho_{j}\right)$, and, for $i\geq 1$, it yields $\sum_{j=2}^{df-1}\rho_{j}^{i}(1-\nu_{j})\lambda_{j}^{2}+(1-\nu_{df})\lambda_{df}^{2}\geq\frac{\nu_{df}}{i+1}\left(df-2-\sum_{j=2}^{df-1}\rho_{j}^{i+1}\right)$. Inequalities (6) follow.
Proofs of Lemmas 1-4
Proof of Lemma 1: Let $\hat{U}$ be the $T\times k$ matrix whose orthonormal columns are the eigenvectors for the $k$ largest eigenvalues of matrix $\hat{V}_{\varepsilon}^{-1/2}\hat{V}_{y}\hat{V}_{\varepsilon}^{-1/2}$. Those eigenvalues are $1+\hat{\gamma}_{j}$, j=1,,kj=1,...,k, while it holds $\hat{F}=\hat{V}_{\varepsilon}^{1/2}\hat{U}\hat{\Gamma}^{1/2}$, where $\hat{\Gamma}=diag(\hat{\gamma}_{1},...,\hat{\gamma}_{k})$. We have $I_{T}-\hat{U}\hat{U}^{\prime}=I_{T}-\hat{V}_{\varepsilon}^{-1/2}\hat{F}\hat{\Gamma}^{-1}\hat{F}^{\prime}\hat{V}_{\varepsilon}^{-1/2}$ $I_{T}-\hat{V}_{\varepsilon}^{-1/2}\hat{F}(\hat{F}^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{F})^{-1}\hat{F}^{\prime}\hat{V}_{\varepsilon}^{-1/2} =$$\hat{V}_{\varepsilon}^{-1/2}M_{\hat{F},\hat{V}_{\varepsilon}}\hat{V}_{\varepsilon}^{1/2}=\hat{V}_{\varepsilon}^{1/2}M_{\hat{F},\hat{V}_{\varepsilon}}^{\prime}\hat{V}_{\varepsilon}^{-1/2}.$ Thus, $\hat{S}=(I_{T}-\hat{U}\hat{U}^{\prime})\left(\hat{V}_{\varepsilon}^{-1/2}\hat{V}_{y}\hat{V}_{\varepsilon}^{-1/2}-I_{T}\right)(I_{T}-\hat{U}\hat{U}^{\prime}).$ By the spectral decomposition of $\hat{V}_{\varepsilon}^{-1/2}\hat{V}_{y}\hat{V}_{\varepsilon}^{-1/2}$, we get $(I_{T}-\hat{U}\hat{U}^{\prime})\left(\hat{V}_{\varepsilon}^{-1/2}\hat{V}_{y}\hat{V}_{\varepsilon}^{-1/2}-I_{T}\right)$ $(I_{T}-\hat{U}\hat{U}^{\prime})=\sum_{j=k+1}^{T}\hat{\gamma}_{j}\hat{P}_{j}$, where the $\hat{P}_{j}$ are the orthogonal projection matrices onto the eigenspaces for the TkT-k smallest eigenvalues. Then, Part (a) follows. Part (b) is a consequence of the squared Frobenius norm of a symmetric matrix being equal to the sum of its squared eigenvalues. For part (c), let $P_{\hat{F},{\hat{V}}_{\varepsilon}}=I_{T}-M_{\hat{F},\hat{V}_{\varepsilon}}$ and note that $\hat{F}\hat{F}^{\prime}=P_{\hat{F},{\hat{V}}_{\varepsilon}}(\hat{V}_{y}-\hat{V}_{\varepsilon})+(\hat{V}_{y}-\hat{V}_{\varepsilon})P_{\hat{F},{\hat{V}}_{\varepsilon}}^{\prime}-P_{\hat{F},{\hat{V}}_{\varepsilon}}(\hat{V}_{y}-\hat{V}_{\varepsilon})P_{\hat{F},{\hat{V}}_{\varepsilon}}^{\prime}$ $\hat{V}_{y}-\hat{V}_{\varepsilon}-M_{\hat{F},\hat{V}_{\varepsilon}}(\hat{V}_{y}-\hat{V}_{\varepsilon})M_{\hat{F},\hat{V}_{\varepsilon}}^{\prime},$ where the first equality is because the three terms on the RHS are all equal to $\hat{F}\hat{F}^{\prime}$ by (FA2). The conclusion follows from (FA1) and $\hat{V}_{\varepsilon}$ being diagonal.
@@
Proof of Lemma 2: We have $\mathscr{Z}_{n}=\frac{1}{\sqrt{n}}(W\Sigma W^{\prime}-Tr(\Sigma)I_{T})$. Hence, $(\mathscr{Z}_{n})_{tt}=\frac{1}{\sqrt{n}}\sum_{i,j}(w_{i,t}w_{j,t}-1_{\{i=j\}})\sigma_{ij}=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}\zeta^{tt}_{m,n},$ with $\zeta^{tt}_{m,n}=\sum_{i\in I_{m}}[w^{2}_{i,t}-1]\sigma_{ii}+2\sum_{\underset{i<j}{i,j\in I_{m}}}w_{i,t}w_{j,t}\sigma_{ij},$ together with $(\mathscr{Z}_{n})_{ts}=\frac{1}{\sqrt{n}}\sum_{i,j}w_{i,t}w_{j,s}\sigma_{ij}=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}\zeta^{ts}_{m,n},t\neq s,$ with $\zeta^{ts}_{m,n}=\sum_{i\in I_{m}}w_{i,t}w_{i,s}\sigma_{ii}+\sum_{\underset{i<j}{i,j\in I_{m}}}w_{i,t}w_{j,s}\sigma_{ij}+\sum_{\underset{i>j}{i,j\in I_{m}}}w_{i,t}w_{j,s}\sigma_{ij},t\neq s,$ so that $vech(\mathscr{Z}_{n})=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}vech(\zeta_{m,n}),$ where $\zeta_{m,n}$ is the $T\times T$ matrix having element $\zeta^{ts}_{m,n}$ in position $(t,s)$. Hence, $vech(\mathscr{Z}_{n})$ is the row sum of a triangular array $\{vech(\zeta_{m,n})\}_{1\leq m\leq n}$ of independent centered random vectors. Let $\Omega_{m,n}:=V[vech(\zeta_{m,n})]$. Using Assumption 2 (a), we compute (i) $E[(\zeta^{tt}_{m,n})^{2}]=\sum_{i\in I_{m}}(E[w^{4}_{i,t}]-1)\sigma^{2}_{ii}+2\sum_{\underset{i\neq j}{i,j\in I_{m}}}\sigma_{ij}^{2}$; (ii) $E[(\zeta^{ts}_{m,n})^{2}]=\sum_{i\in I_{m}}E[w^{2}_{i,t}w^{2}_{i,s}]\sigma_{ii}^{2}+\sum_{\underset{i\neq j}{i,j\in I_{m}}}\sigma_{ij}^{2},t\neq s$; (iii) $E[\zeta^{tt}_{m,n}\zeta^{ss}_{m,n}]=\sum_{i\in I_{m}}E[w^{2}_{i,t}w^{2}_{i,s}-1]\sigma^{2}_{ii},t\neq s$; (iv) $E[\zeta^{tt}_{m,n}\zeta^{rp}_{m,n}]=\sum_{i\in I_{m}}E[w^{2}_{i,t}w_{i,r}w_{ip}]\sigma^{2}_{ii},r\neq p$; (v) $E[\zeta^{ts}_{m,n}\zeta^{rp}_{m,n}]$ $\sum_{i\in I_{m}}E[w_{i,t}w_{i,s}w_{i,r}w_{i,p}]\sigma^{2}_{ii},t\neq s,r\neq p.$ It follows that $V[vech(\mathscr{Z}_{n})]=\frac{1}{n}\sum_{m=1}^{J_{n}}\Omega_{m,n}=D_{n}+\kappa_{n}I_{\frac{T(T+1)}{2}}=\Omega_{n}$. The eigenvalues of $D_{n}$ are bounded away from 0 under Assumption 8 (b), because for any unit vector $\xi\in\mathbb{R}^{T(T+1)/2}$, we have $\xi^{\prime}D_{n}\xi\geq\frac{1}{n}\sum_{i=1}^{n}1_{i\in\bar{S}}\sigma_{ii}^{2}\xi^{\prime}V[vech(w_{i}w_{i}^{\prime})]\xi\geq\underline{c}\frac{1}{n}\sum_{i=1}^{n}1_{i\in\bar{S}}\sigma_{ii}^{2}\geq\underline{c}\left(1-\frac{1}{n}\sum_{i=1}^{n}(1-1_{i\in\bar{S}})\sigma_{ii}\right)^{2}\geq\underline{c}\left(1-\bar{C}\frac{1}{n}\sum_{i=1}^{n}(1-1_{i\in\bar{S}})\right)^{2}\geq\frac{\underline{c}}{4}$, for all $n$. We use the multivariate Lyapunov condition $\|\Omega^{-1/2}_{n}\|^{4}\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[\|vech(\zeta_{m,n})\|^{4}]\to 0$ to invoke a CLT. Since $\|A^{-1/2}\|^{4}\leq\frac{k^{2}}{\delta^{2}_{k}(A)}$ and $\|x\|^{4}\leq k\sum_{j=1}^{k}x^{4}_{j}$, for any $k\times k$ positive semi-definite matrix $A$ and $k\times 1$ vector $x$, it suffices to check that $\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[(\zeta^{ts}_{m,n})^{4}]\to 0,$ for all $t,s$. Besides, we can show that there exists a constant $M>0$, such that $E[(\zeta^{ts}_{m,n})^{4}]\leq Mb_{m,n}^{2(1+\delta)},$ for all $m,n,t,s$. We get $\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}E[(\zeta^{ts}_{m,n})^{4}]\leq M\frac{1}{n^{2}}\sum_{m=1}^{J_{n}}b_{m,n}^{2(1+\delta)}=Mn^{2\delta}\sum_{m=1}^{J_{n}}B_{m,n}^{2(1+\delta)}=o(1)$, under Assumption 2 (d). Then, $\Omega_{n}^{-1/2}vech(\mathscr{Z}_{n})\Rightarrow N(0,I_{\frac{T(T+1)}{2}})$ by the multivariate Lyapunov CLT. Under Assumptions 8 (a)-(c), $\Omega_{n}\rightarrow\Omega$ follows from the Slutsky theorem, and $\Omega$ is positive definite.
@@
$\psi^{(l)}(0)=\frac{(l-1)!}{2}\sum_{j=1}^{df}\nu_{j}^{l-1}\left[\nu_{j}+l(1-\nu_{j})\lambda_{j}^{2}\right],\quad l\geq 0.$ (A.6)
By using the Faa di Bruno formula for the derivatives of a composite function, we have $\frac{d^{l}}{du^{l}}e^{\psi(u)}=e^{\psi(u)}B_{l}(\psi^{\prime}(u),\psi^{{}^{\prime\prime}}(u),...,\psi^{(l)}(u))$, where $B_{l}$ is the $l$th complete exponential Bell’s polynomial (Bell (1934)). Hence, $\Psi^{(l)}(0)=B_{l}(\psi^{\prime}(0),\psi^{{}^{\prime\prime}}(0),...,\psi^{(l)}(0))$. The complete Bell’s polynomials satisfy the recurrence relation $B_{l+1}(x_{1},x_{2},...,x_{l+1})=\sum_{i=0}^{l}\binom{l}{i}B_{l-i}(x_{1},...,x_{l-i})x_{i+1}$. Thus, $\Psi^{(l+1)}(0)=\sum_{i=0}^{l}\binom{l}{i}\Psi^{(l-i)}(0)\psi^{(i+1)}(0)$. After standardization with the factorial term, and using equation (A.6), the conclusion follows.
@@
B Additional theory
B.1 Pseudo likelihood and PML estimator
The FA estimator is the PML estimator based on the Gaussian likelihood function obtained from the pseudo model $y_{i}=\mu+F\beta_{i}+\varepsilon_{i}$ with $\beta_{i}\sim N(0,I_{k})$ and $\varepsilon_{i}\sim N(0,V_{\varepsilon})$ mutually independent and i.i.d. across i=1,,ni=1,...,n. Then, $y_{i}\sim N(\mu,\Sigma(\theta))$ under this pseudo model, where $\Sigma(\theta):=FF^{\prime}+V_{\varepsilon}$ and $\theta:=(vec(F)^{\prime},diag(V_{\varepsilon})^{\prime})^{\prime}\in\mathbb{R}^{r}$ with $r=(k+1)T$. It yields the pseudo log-likelihood function $\hat{L}(\theta,\mu)=-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2n}\sum_{i=1}^{n}(y_{i}-\mu)^{\prime}\Sigma(\theta)^{-1}(y_{i}-\mu)=-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2}Tr\left(\hat{V}_{y}\Sigma(\theta)^{-1}\right)-\frac{1}{2}(\bar{y}-\mu)^{\prime}\Sigma(\theta)^{-1}(\bar{y}-\mu),$ up to constants, where $\bar{y}=\frac{1}{n}\sum_{i=1}^{n}y_{i}$ and $\hat{V}_{y}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\bar{y})(y_{i}-\bar{y})^{\prime}$. We concentrate out parameter $\mu$ to get its estimator $\hat{\mu}=\bar{y}$. Then, estimator $\hat{\theta}=(vec(\hat{F})^{\prime},diag(\hat{V}_{\varepsilon})^{\prime})^{\prime}$ is defined by the maximization of
$\hat{L}(\theta):=-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2}Tr\left(\hat{V}_{y}\Sigma(\theta)^{-1}\right),$ (B.1)